The population of a place increased to 54000 in 2003 at a rate of 5% per annum 
what would be its population in 2005? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

It is given that, population in the year 2003 = 54,000
(Population in 2005)=54000(1 + 5/100 )^2
= 54000(1 +1/20 )^2
= 54000(1 + 21/20 )^2
= 59535
∴ The population in the year 2005 would be 59,535.

​Thanks